Reports 1-1 of 1 Clear search Modify search
Virgo Runs (O4b)
irace - 16:07 Tuesday 07 May 2024 (64186) Print this report
Operator Report - Morning shift

Upon arrival, I found the ITF locked and in SCIENCE mode.

At 6:05 UTC I set the MAINTENANCE mode. Here is the list of the activities communicated to the Control Room:

- Cleaning of the Central Building and the DET Lab (external firm managed by Menzione);

- TCS chiller refill (Menzione);

- Replacement of the electric post-heating coil for the AHU 1_SX of the Mode Cleaner (external firm managed by Soldani);

- Checks on the HVAC plants (Soldani);

- ELE_QGBT process update (Ballardin);

- ENV_MCB_MAG_N magnetometer substitution, see #64189 (Paoletti)

During the shift, I performed these periodic tasks:

- At 6:10 UTC I set the SINGLE_BOUNCE_NI state. OMC locked from 6:30 UTC to 6:35 UTC. OMC scan performed from 6:45 UTC to 7:06 UTC;

- Check of IMC working point and position from 7:20 UTC to 7:30 UTC;

- Acquisition of TCS thermal camera references at 7:47 UTC;

- At 9:00 UTC checks on TCS powers (see table below for results);

WI pickoff 0.271 0 0.212
NI Pickoff 0.662 0.053 0.590

At 9:03 UTC the SNEB horizontal position control loop opened. I closed it at 9:28 UTC.

NCAL flag on DMS became red at 8:20 UTC. Experts informed.

As requested by Degallaix, I performed a slow FSR scan with the green laser from 10:44 UTC to 11:13 UTC. After that I set the PREPARE_SCIENCE mode.

During the lock acquisition a CITF alignment was required. The SUS_PR and DRMI_LOCK Metatron nodes got stalled, requiring a restart of the latter.

I left the ITF in LOW_NOISE_3_SQZ state and SCIENCE mode at 13:10 UTC.


Images attached to this report
Search Help